# Optimal Client Sampling in Federated Learning with Client-Level Heterogeneous Differential Privacy

### v1

@@ 4 Sparsification-Amplified GDPFed with Optimal Client Sampling
Given a vector $x\in\mathbb{R}^{d}$ and a sparsification parameter $k\in[d]$. The $\mathrm{Top}_{k}(\cdot)$ holds that $\mathbb{E}\|\mathrm{Top}_{k}(x)-x\|^{2}\leq\phi\|x\|^{2}$, where $\phi$ is a sparsification error coefficient.
It is evident that a smaller $k$ results in a larger $\phi$, thereby leading to a greater sparsification error. Therefore, $k_{m}$ should be carefully selected in order to successfully leverage its benefit. In the literature, $\phi$ is typically set to $1-k/d$ [25, 57] or $(1-k/d)^{2}$ [49] to measure the sparsification error. In this work, we choose $\phi=(1-k/d)^{2}$ as it provides a tighter bound.
Convergence Analysis of GDPFed. We now present a detailed convergence analysis of the sparsification-amplified GDPFed. In the following, we present several important assumptions that help us conduct the convergence analysis.
@@
The local objective $f_{m,i}(\cdot)$ of each client $i\in\mathcal{G}_{m}$ in any group $m\in[M]$, is $L$-smooth with constant $L>0$; i.e., for all $x,y\in\mathbb{R}^{d}$, $\left\|\nabla f_{m,i}(x)-\nabla f_{m,i}(y)\right\|\leq L\left\|x-y\right\|,$ which implies $f_{m,i}(x)-f_{m,i}(y)\leq\nabla f_{m,i}(x)^{\top}(y-x)+(L/2)\left\|x-y\right\|^{2}$.
@@
There exist two constants $\beta^{2}\geq 1$ and $\kappa^{2}\geq 0$ such that $\sum_{m=1}^{M}\omega_{m}\sum_{i\in\mathcal{G}_{m}}\left\|\nabla f_{m,i}(x)\right\|^{2}\leq\beta^{2}\|\sum_{m=1}^{M}\omega_{m}\sum_{i\in\mathcal{G}_{m}}\nabla f_{m,i}(x)\|^{2}+\kappa^{2}.$ If all local objective functions are identical, the inequality holds with $\beta^{2}=1$ and $\kappa^{2}=0$.
Note that Assumption 1 and Assumption 2 are commonly used in the theoretical analysis of distributed learning systems [57, 44, 25]. In particular, Assumption 2 bounds the coordinate-wise variance of local gradients [24]. Meanwhile, Assumption 3 captures inter-client heterogeneity in FL [15, 4, 28]. With the above assumptions, we provide the convergence result of GDPFed under the general non-convex setting in Theorem 3.
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f(\theta^{t})\right\|^{2}$ $\leq\frac{8\left(f(\theta^{0})-f^{*}\right)}{\eta T\tau}+\mu_{1}\kappa^{2}+\mu_{2}\sum_{m=1}^{M}\omega_{m}(\phi_{m}+1)d\zeta_{m}^{2}+\mu_{3}\sum_{m=1}^{M}\frac{k_{m}\omega_{m}^{2}C^{2}\sigma_{m}^{2}}{r_{m}q_{m}},$
@@
The detailed proof is given in Appendix D. ∎
@@
If $\phi_{m}=0,\ \forall m\in[M]$, meaning no sparsification is applied, the first three terms on the right-hand side of the convergence bound correspond to the optimization error of FedAvg. In particular, the third term captures group-wise heterogeneity in model updates, which are influenced by the group-wise sparsification parameters $k_{m}$. Specifically, applying more aggressive sparsification (i.e., smaller $k_{m}$) increases the heterogeneity among per-group model updates. However, as reflected in the final term of the bound, a smaller $k_{m}$ reduces the privacy error introduced by DP, confirming our analysis in Section 3. This highlights a fundamental trade-off: selecting an appropriate $k_{m}$ is crucial for balancing sparsification and privacy errors, thereby minimizing the overall convergence error. Hence, by directly minimizing the errors in the third and last terms that are related to $k_{m}$, we obtain we obtain a coarse closed-form expression for the optimal sparsification level for the group $m$: $k^{*}_{m}/d=1-2\omega_{m}\sigma_{m}^{2}/(\eta\tau\mu_{4}r_{m}^{2}),\ \forall m\in[M],$ where $\mu_{4}=32\eta\tau+\eta+\eta/\tau$. At this case, $\phi^{*}_{m}$ is given by $\phi^{*}_{m}=4\omega_{m}^{2}\sigma_{m}^{4}/(\eta\tau\mu_{4}r_{m}^{2})^{2}$ (see the sketch of the derivation in Appendix E). This yields a tighter upper bound for the convergence error.
Optimal Client Sampling Ratios. Building on the convergence analysis of sparsification-amplified GDPFed, we now discuss how to determine the optimal client sampling ratio for each group. To ensure that the global model trained by GDPFed converges to a better optimum, it is desirable to minimize the true gradient of the objective function (i.e., the left-hand side of the convergence result). However, directly minimizing this function is typically infeasible in practice, as $\nabla f(\theta^{t})$ is a high-dimensional, non-convex function. An alternative approach is optimizing its upper bound (i.e., the right-hand side of the convergence bound), which approximates optimizing the objective function. Notably, only the third and last terms in the bound are influenced by the client sampling ratios. This leads to the constrained minimization problem formulated in Problem 1.
@@
s.t. $r_{m}=q_{m}|\mathcal{G}_{m}|,\ \sum_{m\in[M]}r_{m}=qn,$
where $\mu_{4}=32\eta\tau+\eta+\eta/\tau$ and $\mu_{5}=4/(\eta\tau)$. The formulation sketch is provided in Appendix F.
@@
The optimal sparsification error coefficient $\phi^{*}_{m}$ is defined as given in Remark 3. If there is no sparsification applied, then $\phi^{*}_{m}=0,\ \forall m\in[M]$. The noise multiplier $\sigma_{m}^{2}$ required to satisfy the $(\epsilon_{m},\delta)$-DP for group $m$ is derived in Theorem 1. All parameters in Problem 1 are now the settings of the system, except for the decision variables. Therefore, the optimal client sampling ratios for each group in GDPFed can be efficiently obtained by solving this minimization problem. As Problem 1 is a non-convex optimization problem, one can resort to existing solvers in practice, such as optimization libraries in Python (e.g., scikit-learn [46]), to obtain a feasible solution.
@@
With the optimal client sampling ratios derived from solving Problem 1, which minimizes the convergence upper bound in Theorem 3, GDPFed+ converges to a better minimum than GDPFed, thereby enhancing model utility. Importantly, GDPFed+ still satisfies the per-group privacy guarantees in Theorem 1, the overall privacy guarantee in Theorem 2, and the convergence bound in Theorem 3.
5 Empirical Evaluation
Table 1: Configurations for each dataset.
[…]
FMNIST $6,000$ $50$ $1.5$ $(0.5,1.5,3.0)$ $(0.69,1.89,3.42)$-$2$
SVHN $6,000$ $100$ $1.0$ $(0.5,1.5,3.0)$ $(1.66,4.67,8.68)$-$5$
Shakespeare $714$ $50$ $1.0$ $(0.5,1.5,3.0)$ $(4.79,9.83,15.21)$-$10$
CIFAR-10 $600$ $100$ $1.5$ $(2.0,6.0,12.0)$ $(3.61,9.62,16.77)$-$10$
@@ 5 Empirical Evaluation
Datasets and Settings. Our evaluation covers four benchmark datasets for DPFL: Fashion MNIST (FMNIST) [55], SVHN [45], CIFAR-10 [30], and Shakespeare [10]. Correspondingly, we adopt a $2$-layer CNN for FMNIST, a $3$-layer CNN for SVHN, a ResNet-18 [21] for CIFAR-10, and an LSTM model for Shakespeare. We conduct experiments in cross-device FL settings with $n$ clients. Datasets, excluding Shakespeare, are evenly partitioned (i.e., IID) across clients; Shakespeare is used in its natural non-IID form. Following prior works [3, 7, 29] that simulate heterogeneous privacy requirements, clients are assigned to one of three groups, each associated with a distinct minimum privacy budget $(\epsilon_{1},\epsilon_{2},\epsilon_{3})$. By default, clients are evenly distributed among three groups. For GDPFed+, the optimal client sampling ratios $(q_{1},q_{2},q_{3})$ derived by solving Problem 1. By default, the sparsification levels $k_{m}/d$ for each group are $(0.7,0.8,0.9)$. We provide detailed empirical observations on the effects of varying sparsification levels and offer practical suggestions for selecting appropriate sparsification configurations in Appendix B.4. We summarize the system configurations in Table 1. Detailed experimental settings are provided in Appendix B.2. All experiments are repeated $3$ times with different seeds.
[…]
@@
Baselines. We compare against four baselines to demonstrate the effectiveness of GDPFed+. Specifically, we include two important baselines: Pure FedAvg (P-FedAvg), a non-private FL case that represents the upper bound of model utility, and client-level DP-FedAvg (DP-FedAvg), which enforces the strictest privacy requirement across all clients. Moreover, our comparisons include IDP-FedAvg [7] and PFA [32] (see the rationale behind selecting these methods in Appendix B.2).
@@
Experimental Results. We begin by presenting the convergence curves of representative methods in Figure 1 and the corresponding test accuracies in Table 2. As shown in Figure 1, DP-FedAvg suffers from substantial performance degradation compared to P-FedAvg, due to applying the strictest privacy level uniformly across all clients. In contrast, GDPFed, which enforces group-level DP guarantees, consistently converges to better optima across all datasets by reducing the noise magnitude for clients with looser privacy requirements. Building upon this, GDPFed+ further integrates model sparsification with optimal per-group client sampling ratios, resulting in enhanced performance and achieving the highest test accuracies on all datasets. Moreover, GDPFed+ exhibits a more stable convergence process compared to other baselines. As shown in Table 2, GDPFed+ outperforms two competitive baselines, PFA and IDP-FedAvg, with average accuracy gains of $+3.17\%$ and $+2.93\%$, respectively. These results demonstrate the empirical effectiveness of GDPFed+ in enhancing model utility while preserving client-level HDP guarantees.
@@
Table 3: Detailed noise multipliers and $\Lambda$ for different methods.
Method $\sigma^{2}$ /// $(\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{3})$ $\Lambda$
DP-FedAvg $2.26$ $5.09$
GDPFed $(2.26,0.90,0.53)$ $2.77$
GDPFed-$\mathrm{op}$ $(1.42,0.87,0.70)$ $0.57$
GDPFed+ $(\mathbf{1.42},\mathbf{0.87},\mathbf{0.70})$ $\mathbf{0.49}$

### v3

@@ V-B Bounded Sparsification Error
Given a vector $x\in\mathbb{R}^{d}$ and a sparsification parameter $k\in[0,d]$. We have $\mathbb{E}\|\mathrm{Top}_{k}(x)-x\|^{2}\leq\phi\|x\|^{2}$, where $\phi$ is a sparsification error coefficient.
It is evident that a smaller $k$ results in a larger $\phi$, thereby leading to a greater sparsification error. Therefore, $k_{m}$ should be carefully selected in order to successfully leverage its benefit. In the literature, $\phi$ is typically set to 1k/d1-k/d [25, 58] or $(1-k/d)^{2}$ [50] to measure the sparsification error. In this work, we choose $\phi=(1-k/d)^{2}$ as it provides a tighter bound.
V-C Convergence Analysis of GDPFed
We first present several important assumptions that help us conduct the convergence analysis.
@@ V-C Convergence Analysis of GDPFed
The local objective $f_{m,i}(\cdot)$ of each client $i\in\mathcal{G}_{m}$ in any group $m\in[M]$, is $L$-smooth with constant $L>0$; i.e., for all $x,y\in\mathbb{R}^{d}$, $\left\|\nabla f_{m,i}(x)-\nabla f_{m,i}(y)\right\|\leq L\left\|x-y\right\|$.
@@
There exist $\beta^{2}\geq 1$ and $\kappa^{2}\geq 0$ such that $\sum_{m=1}^{M}\omega_{m}\sum_{i\in\mathcal{G}_{m}}\left\|\nabla f_{m,i}(x)\right\|^{2}\leq\beta^{2}\|\sum_{m=1}^{M}\omega_{m}\sum_{i\in\mathcal{G}_{m}}\nabla f_{m,i}(x)\|^{2}+\kappa^{2}.$ With identical local objectives, the inequality holds with $\beta^{2}=1$ and $\kappa^{2}=0$.
Note that Assumption 12 are commonly used in the theoretical analysis of distributed learning systems [57, 44, 25]. In particular, Assumption 2 bounds the coordinate-wise variance of local gradients [24]. Meanwhile, Assumption 3 captures inter-client heterogeneity in FL [14, 3, 28]. With the above assumptions, we provide the convergence result of GDPFed under the general non-convex setting in Theorem 2.
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\nabla f(\theta^{t})\right\|^{2}$ $\leq\frac{8\left(f(\theta^{0})-f^{*}\right)}{\eta T\tau}+\mu_{1}\kappa^{2}$
$+\mu_{2}\sum_{m=1}^{M}\omega_{m}(\phi_{m}+1)d\zeta_{m}^{2}$
$+\mu_{3}\sum_{m=1}^{M}\frac{k_{m}\omega_{m}^{2}C^{2}\sigma_{m}^{2}}{r_{m}q_{m}},$
@@
The detailed proof is given in Appendix G. ∎
@@
If $\phi_{m}=0,\ \forall m\in[M]$, meaning no sparsification is applied, the first three terms on the right-hand side of the convergence bound correspond to the optimization error of FedAvg. In particular, the third term captures group-wise heterogeneity in model updates, which are influenced by the group-wise sparsification parameters $k_{m}$. Specifically, applying more aggressive sparsification (i.e., smaller $k_{m}$) increases the heterogeneity among per-group model updates. However, as reflected in the final term of the bound, a smaller $k_{m}$ reduces the privacy error introduced by DP, confirming our analysis in Section IV. This highlights a fundamental trade-off: selecting an appropriate $k_{m}$ is crucial for balancing sparsification and privacy errors, thereby minimizing the overall convergence error. Hence, by directly minimizing the errors in the third and last terms that are related to $k_{m}$, we obtain a coarse closed-form expression for the optimal sparsification level for the group $m$: $k^{*}_{m}/d=1-2\omega_{m}\sigma_{m}^{2}/(\eta\tau\mu_{4}r_{m}^{2}),\ \forall m\in[M],$ where $\mu_{4}=32\eta\tau+\eta+\eta/\tau$. At this case, $\phi^{*}_{m}$ is given by $\phi^{*}_{m}=4\omega_{m}^{2}\sigma_{m}^{4}/(\eta\tau\mu_{4}r_{m}^{2})^{2}$ (see the sketch of the derivation in Appendix H). This yields a tighter upper bound for the convergence error. Importantly, $\phi_{m}^{*}$ can be directly specified from the system configuration.
V-D Optimal Client Sampling Ratios
Building on the convergence analysis of sparsification-amplified GDPFed, we now discuss how to determine the optimal client sampling ratio for each group. To ensure that the global model trained by GDPFed converges to a better optimum, it is desirable to minimize the true gradient of the objective function (i.e., the left-hand side of the convergence result). However, directly minimizing this function is typically infeasible in practice, as $\nabla f(\theta^{t})$ is a high-dimensional, non-convex function. An alternative approach is optimizing its upper bound (i.e., the right-hand side of the convergence bound), which approximates optimizing the objective function. Notably, only the third and last terms in the bound are influenced by the client sampling ratios. This leads to the constrained minimization problem formulated in Problem 1.
@@ V-D Optimal Client Sampling Ratios
s.t. $r_{m}=q_{m}|\mathcal{G}_{m}|,\ \sum_{m\in[M]}r_{m}=qn,\$
where $\mu_{4}=32\eta\tau+\eta+\eta/\tau$ and $\mu_{5}=4/(\eta\tau)$.
@@
The optimal sparsification error coefficient $\phi^{*}_{m}$ is defined as given in Remark 3. If there is no sparsification applied, then $\phi^{*}_{m}=0,\ \forall m\in[M]$. The noise multiplier $\sigma_{m}^{2}$ required to satisfy the $(\epsilon_{m},\delta)$-DP for group $m$ is derived in Theorem 1. All parameters in Problem 1 are now determined by the system settings, as terms such as $L$ and $\zeta$ have been eliminated (see the formulation sketch in Appendix I for details), leaving only the decision variables. Therefore, the optimal client sampling ratios for each group in GDPFed can be efficiently obtained by solving this minimization problem. As Problem 1 is a non-convex optimization problem, one can resort to existing solvers in practice, such as optimization libraries in Python (e.g., scikit-learn [47]), to obtain a feasible solution.
@@
With the optimal client sampling ratios derived from solving Problem 1, which minimizes the convergence upper bound in Theorem 2, GDPFed+ converges to a better minimum than GDPFed, thereby enhancing model utility. Importantly, GDPFed+ still satisfies the per-group privacy guarantees in Theorem 1, the overall privacy guarantee in Corollary 1, and the convergence bound in Theorem 2. In the following evaluation section, we empirically validate the performance of the proposed GDPFed and GDPFed+ methods, which are equipped with per-group sparsification and optimal client sampling ratios, against several baseline approaches.
VI Empirical Evaluation
VI-A Experimental Settings
TABLE I: Detailed system configurations for each dataset.
Dataset $n$ $T$ $C$ $(\epsilon_{1},\epsilon_{2},\epsilon_{3})$ $(q_{1},q_{2},q_{3})$-$q$ $(\%)$
FMNIST 6,0006{,}000 $50$ $1.5$ $(0.5,1.5,3.0)$ (0.69,1.89,3.42)(0.69,1.89,3.42)-22
SVHN 6,0006{,}000 $100$ $1.0$ $(0.5,1.5,3.0)$ (1.66,4.67,8.68)(1.66,4.67,8.68)-55
Shakespeare $714$ $50$ $1.0$ $(0.5,1.5,3.0)$ (4.79,9.83,15.21)(4.79,9.83,15.21)-$10$
CIFAR-10 $600$ $100$ $1.5$ $(2.0,6.0,12.0)$ (3.61,9.62,16.77)(3.61,9.62,16.77)-$10$
VI-A1 Dataset settings
Our evaluation covers four benchmark datasets for HDPFL: Fashion MNIST (FMNIST) [56], SVHN [45], CIFAR-10 [30], and Shakespeare [9]. Correspondingly, we adopt a $2$-layer CNN for FMNIST, a $3$-layer CNN for SVHN, a ResNet-18 [21] for CIFAR-10, and an LSTM model for Shakespeare. We conduct experiments in cross-device FL settings with $n$ clients. Client local datasets are assumed to be independent and identically distributed (IID), except Shakespeare, which is evaluated in its inherent non-IID form. We summarize the system configurations in Table I.
VI-A2 Baselines
We compare against four baselines to demonstrate the effectiveness of GDPFed+. Specifically, we include two important baselines: Pure FedAvg (P-FedAvg), a non-private FL case that serves as a strong reference point for model utility without privacy noise, and client-level DP-FedAvg (DP-FedAvg), which enforces the strictest privacy requirement across all clients. Moreover, our comparisons include state-of-the-art methods IDP-FedAvg [6] and PFA [32]. We follow their default settings, with further details provided in Appendix C.
VI-A3 Training settings
In all experiments, local clients use stochastic gradient descent (SGD) as the optimizer with a learning rate of $\eta=0.1$ and a decay ratio of $0.99$. For the FMNIST, SVHN, CIFAR-10, and Shakespeare datasets, we set the momentum coefficient to $0.0$, $0.0$, $0.5$, and $0.9$, the number of local training iterations $\tau$ to $5$, $25$, 55, and $30$, and the batch size to $10$, $10$, $50$, and $4$, respectively. We set the uniform DP failure parameter as $\delta=1/n^{1.1}$, following the recommendation in [25, 38]. To reweight the per-group model updates, we set the reweighting parameter $\omega_{m}=(1/qn)\cdot r_{m}^{2}/\sum_{m\in[M]}r_{m}^{2}$ for all $m\in[M]$. This reweighting strategy prioritizes groups with higher expected client participation and helps reduce the total noise added to the aggregated model. Note that to ensure a fair comparison, we adopt this reweighting parameter for all group-based methods, including PFA, IDP-FedAvg, GDPFed, and GDPFed+. Additional baseline settings are provided in Appendix C. All experiments are repeated three times with different random seeds to ensure statistical reliability.
VI-A4 System settings
@@
Following prior works [2, 6, 29] that simulate heterogeneous privacy requirements, clients are assigned to one of three groups, each associated with a distinct minimum privacy budget $(\epsilon_{1},\epsilon_{2},\epsilon_{3})$. By default, clients are evenly distributed among three groups. For GDPFed+, the optimal client sampling ratios $(q_{1},q_{2},q_{3})$ are derived by solving Problem 1. By default, the sparsification levels $k_{m}/d$ for each group are $(0.7,0.8,0.9)$.
VI-A5 Hardware settings
All experiments were conducted on a Linux-based internal compute cluster equipped with $8$ NVIDIA RTX A60006000 GPUs (each with $48$ GB of memory) and AMD EPYC $7763$ $64$-core CPUs.
VI-B Experimental Results
Figure 1: Optimal client sampling ratios under varying $\epsilon_{1}$. Sampling ratio for each group is adjusted dynamically to satisfy the global constraint with fixed $\epsilon_{2}$ and $\epsilon_{3}$.
VI-B1 Dynamics of client sampling ratios
We first demonstrate how the optimal client sampling ratios dynamically adjust in response to varying privacy budgets, as dictated by our optimization formulation in Problem 1. Specifically, we conduct experiments on the FMNIST and CIFAR-10 datasets. In both cases, we fix the privacy budgets for Group $2$ and Group $3$ as $\epsilon_{2}=1.5$ and $\epsilon_{3}=3.0$ for FMNIST and $\epsilon_{2}=6.0$ and $\epsilon_{3}=12.0$ for CIFAR-10, respectively, and vary $\epsilon_{1}$ to examine how the optimal sampling ratios evolve. The resulting trends are illustrated in Figure 1. Across both datasets, we observe a consistent pattern governed by the optimization objective. When $\epsilon_{1}$ is small, the corresponding sampling ratio $q_{1}$ decreases to accommodate the stronger noise required for stricter privacy. As $\epsilon_{1}$ increases, $q_{1}$ rises accordingly, while $q_{2}$ and $q_{3}$ adjust downward to maintain the global constraint $\sum_{m\in[M]}r_{m}=qn$. Notably, when $\epsilon_{1}=\epsilon_{2}=1.5$ for FMNIST and $\epsilon_{1}=\epsilon_{2}=6.0$ for CIFAR-10, Groups $1$ and $2$ yield identical sampling ratios.
Figure 2: Convergence curve comparison between P-FedAvg, DP-FedAvg, GDPFed, and GDPFed+.
TABLE II: Test accuracy ($\%$) of baselines, GDPFed, and GDPFed+ on each dataset.
@@ VI-B1 Dynamics of client sampling ratios
VI-B2 Convergence and main results
We present the convergence curves of P-FedAvg, DP-FedAvg, GDPFed, and GDPFed+ in Figure 2, together with the test accuracies in Table II. As shown in Figure 2, DP-FedAvg exhibits clear performance degradation compared to P-FedAvg. This is because DP-FedAvg enforces the strictest privacy budget uniformly across clients, resulting in strong noise injection at every training round. On the Shakespeare task, DP-FedAvg initially achieves a modest accuracy improvement, but its performance subsequently degrades over time. In the early training rounds, the cumulative DP noise remains limited, allowing the model to capture useful patterns. However, as training progresses, the accumulated noise grows across rounds, and due to the high sensitivity of the Shakespeare task to DP noise, it gradually overwhelms the learning signal of the model, ultimately causing model collapse.
@@ VI-B2 Convergence and main results
In contrast, GDPFed enforces DP at the group level rather than using a single global privacy budget. By reducing unnecessary noise for clients with looser privacy requirements, GDPFed consistently converges to better solutions across all datasets. Building upon this, GDPFed+ further integrates model sparsification with optimal per-group client sampling ratios, resulting in enhanced performance and achieving the highest test accuracies on all datasets. Moreover, GDPFed+ exhibits a more stable convergence process compared to other baselines. As shown in Table II, GDPFed+ achieves an average accuracy of $56.93\%$ across all datasets, representing a $5.75\%$ improvement over GDPFed’s accuracy of $51.18\%$. Moreover, GDPFed+ surpasses two state-of-the-art baselines, PFA and IDP-FedAvg, with average accuracy gains of $3.17\%$ and $2.93\%$, respectively. These results empirically demonstrate the effectiveness of GDPFed+ in enhancing model utility while maintaining client-level HDP guarantees.
@@
TABLE III: Noise multipliers ($\sigma^{2}$ or $(\sigma^{2}_{1},\sigma^{2}_{2},\sigma^{2}_{3})$) and noise magnitude ($\Lambda$) on FMNIST and CIFAR-10 across methods.
Method FMNIST CIFAR-10
DP-FedAvg $2.26$ / $5.09$ $3.52$ / $7.91$
GDPFed $(2.26,0.90,0.53)$ / $2.77$ $(3.52,0.95,0.49)$ / $3.72$
GDPFed-$\mathrm{opc}$ $(1.42,0.87,0.70)$ / $0.57$ $(0.98,0.91,0.83)$ / $0.64$
GDPFed+ $(\mathbf{1.42},\mathbf{0.87},\mathbf{0.70})$ / $\mathbf{0.49}$ $(\mathbf{0.98},\mathbf{0.91},\mathbf{0.83})$ / $\mathbf{0.56}$
VI-B3 Noise amount analysis